Line ax+by+p=0 makes an angle 45 degree with xcosA+ysinA=p, p€R+. if+these lines and the line xsinA-ycosA=0 are concurrent, then

(1)asquare+bsquare=1
(2)asquare+bsquare=2
(3)2(asquare+bsquare)=1
(4)none of these

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear Chandrakesh

ax+by+p=0                m1 = -a/b

xcosA+ysinA=p         m2 = -cosA/sinA


 tan45=1= |m1-m2|/|1+m1m2|

     |1+m1m2| = |m1-m2|

   |bsinA +a cosA| = |asinA -bcosA|

 square

  b2 sin2A + a2cos2A +2absinAcosA =a2 sin2A + b2cos2A -2absinAcosA     .............1


given three lines are concurrent

point of intersection of xcosA+ysinA=p  and xsinA-ycosA=0  are

 x=pcosA

 y =psinA

it will satisfied by 3rd line

so  acosA + bSinA =-p

  square

 b2 sin2A + a2cos2A +2absinAcosA  =p ...................2

from equation 1

a2 sin2A + b2cos2A -2absinAcosA =p2 .......................3

add 2 and 3 equation

 a2  + b2  =2p2


Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin